# Instance-Adaptive Keypoint Learning with Local-to-Global Geometric Aggregation for Category-Level Object Pose Estimation

### v3

@@ IV-B1 Results on REAL275
Table I reports the performance of INKL-Pose compared with state-of-the-art methods on the REAL275 dataset [10]. As shown in the table, our method exhibits excellent performance on most metrics. Specifically, INKL-Pose achieves 84.0% on IoU50, 80.6% on IoU75, 63.2% on 5°2cm, 68.3% on 5°5cm, 80.9% on 10°2cm, and 88.0% on 10°5cm, demonstrating its strong capability in both pose and size estimation under real-world conditions. Notably, INKL-Pose operates directly on RGB-D input and does not rely on any pre-defined shape priors. Compared with CatFormer [12], the state-of-the-art prior-based method, INKL-Pose achieves significant improvements 15.5% and 14.6%, under the strict 5°2cm and 5°5cm metrics, respectively. Furthermore, compared to SpherePose [21], the current best-performing shape-prior-free method, INKL-Pose yields performance gains of 5.0% and 0.9% under the strict 5°2cm and 5°5cm metrics, respectively. In addition, compared to our strong baseline AG-Pose [20], INKL-Pose achieves notable improvements of 6.2% and 3.7% on 5°2cm and 5°5cm metrics, respectively, highlighting the effectiveness of our fine-to-global geometric aggregation strategy. Qualitative comparisons in Fig. 4 further validate the superiority of INKL-Pose.
@@ IV-B2 Results on CAMERA25
Table II reports the performance of INKL-Pose compared with state-of-the-art methods on the CAMERA25 dataset. As shown in the table, INKL-Pose achieves outstanding accuracy across nearly all evaluation metrics. Although its IoU50 metric reaches 94.4%, slightly lower than SpherePose [21] (achieving 94.8%), it significantly outperforms all existing approaches on the remaining metrics, especially under the more rigorous pose accuracy thresholds. Specifically, compared to prior-based methods, INKL-Pose maintains a clear advantage without relying on pre-defined shape priors. In addition, when compared with AG-Pose [20], the current state-of-the-art shape-prior-free method, INKL-Pose achieves notable gains of 1.5% and 1.3% under the strict 5°2cm and 5°5cm metrics, respectively. These results, along with the strong performance observed on the REAL275 dataset, provide compelling evidence of the robustness and effectiveness of our proposed local-to-global geometric feature aggregation framework. Furthermore, INKL-Pose exhibits strong model efficiency with 16.7M parameters, ranking among the most lightweight methods evaluated except for AG-Pose [20] and GPV-Pose [35]. It also achieves a fast inference speed of 36 FPS on an NVIDIA RTX 4090D GPU, highlighting its potential for real-time applications.
@@
IoU25 IoU50 $5^{\circ}2\mathrm{cm}$ $5^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$
GPV-Pose[35] 74.9 50.7 3.5 4.6 17.8 22.7
VI-Net[18] 80.7 56.4 8.4 10.3 20.5 29.1
SecondPose[19] 83.7 66.1 11.0 13.4 25.3 35.7
AG-Pose[20] 82.4 66.0 11.5 12.6 37.4 42.5
AG-Pose (DINOv2)[20] 88.1 76.9 21.3 22.1 51.3 54.3
SpherePose[21] 88.8 72.2 19.3 25.9 40.9 55.3
GCE-Pose[37] - 79.2 24.8 25.7 55.4 58.4
Ours 90.8 81.9 25.2 27.4 57.5 61.9
TABLE III: Quantitative comparisons with state-of-the-art methods on the HouseCat6D Dataset[31]. The best results are in bold.
Figure 5: Visualization results of INKL-Pose on the HouseCat6D dataset. Green bounding boxes denote ground truth, while red bounding boxes represent the predictions from our INKL-Pose. The predicted boxes show high consistency with the ground truth.
@@ IV-B3 Results on HouseCat6D
To further validate the generalization capability of the proposed INKL-Pose, we conduct experiments on the more challenging real-world dataset HouseCat6D [31]. As reported in Table III, INKL-Pose achieves state-of-the-art performance across all evaluation metrics. Compared with the current state-of-the-art GCE-Pose [37], our method surpasses it under all evaluation metric, with gains of 2.7% on IoU50, 0.4% on 5°2cm, 1.7% on 5°5cm, 2.1% on 10°2cm, and 3.5% on 10°5cm. Moreover, Fig. 5 presents qualitative results of the predicted poses, further showcasing the strong generalization ability of INKL-Pose in complex real-world scenarios.
@@ IV-B4 Visualization of Predicted Keypoints
Fig. 6 compares the keypoints predicted by our method with those from the baseline method AG-Pose [20] in the NOCS space. To highlight the differences, keypoints are color-coded based on their NOCS errors, linearly mapped from green (error = 0) to red (error = 0.2). As shown in the figure, AG-Pose produces keypoints with higher NOCS errors and uneven surface coverage. Incorporating our keypoint loss functions into AG-Pose leads to moderate improvements in both accuracy and spatial distribution. In contrast, our INKL-Pose achieves the lowest NOCS errors and generates a more uniform and comprehensive keypoint layout that better captures the object’s surface geometry. For example, in the third column showing a mug with a handle, AG-Pose fails to localize the handle, while its variant with our losses partially recovers the region but still exhibits large errors. In comparison, our method accurately delineates the handle contour, resulting in more precise pose estimation.
@@ IV-C Ablation Studies
In this section, we conduct comprehensive ablation studies to evaluate the contributions of key components proposed in INKL-Pose. All experiments are performed on the REAL275 [10] dataset.
@@ IV-C1 Effect of the Number of Keypoints
We begin by evaluating how the number of predicted keypoints influences the overall pose estimation performance. As shown in Table IV, even with only 16 sparse keypoints, INKL-Pose achieves 56.0% accuracy under the 5°2cm metric, which is already comparable to many existing state-of-the-art methods. As the number of keypoints increases, the performance steadily improves due to a stronger capacity for capturing fine-grained geometric details. The best result is achieved with 96 keypoints, reaching 63.2% on 5°2cm, which is adopted as the default setting in our framework.
$N_{kpt}$ $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
16 56.0 64.3 75.3 85.5
32 57.6 66.0 76.1 87.1
64 59.6 66.7 78.4 87.4
96 63.2 68.3 80.9 88.0
128 59.3 67.7 77.3 87.2
TABLE IV: Effect of the number of keypoints. Default configurations are colored in gray, with best results highlighted in bold.
Method $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
w/ ($\mathcal{L}_{sep}$+$\mathcal{L}_{surf}$) 63.2 68.3 80.9 88.0
w/ ($\mathcal{L}_{div}$+$\mathcal{L}_{ocd}$) 60.9 67.0 79.9 87.2
@@
TABLE V: Effect of keypoint supervision. Default configurations are colored in gray, with best results highlighted in bold.
IV-C2 Effect of Keypoint Supervision
In the Instance-Adaptive Keypoint Detector (IAKD) module (Section III-C), we introduce two auxiliary loss terms, $\mathcal{L}_{surf}$ and $\mathcal{L}_{sep}$, which encourage uniform surface coverage and spatial diversity among keypoints. Table V summarizes the performance under different supervision strategies. As the table shows, without any keypoint supervision (“None”), the model performs poorly across all metrics, highlighting the necessity of proper guidance. When incorporating the surface coverage loss $\mathcal{L}_{ocd}$ proposed in AG-Pose, the performance further degrades, indicating that this loss is not well suited to our framework. In contrast, the proposed surface loss $\mathcal{L}_{surf}$ significantly improves performance over the unsupervised baseline.
IV-C3 Effect of Positional Embeddings
We further analyze the contribution of positional information to feature representation. In INKL-Pose, positional encodings are introduced at two distinct stages: the embedding $\mathcal{E}_{\mathcal{P}}$ in the Pointwise Feature Encoder (Section III-B), which compensates for the loss of absolute position cues, and $\mathcal{E}_{kpt}$ in the LKFA module (Section III-D), which enhances local spatial awareness. As reported in Table VI, removing either embedding individually leads to a noticeable performance drop across all metrics, highlighting their complementary roles in capturing geometric information.
Method $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
@@ IV-C3 Effect of Positional Embeddings
TABLE VI: Effect of positional embeddings. Default configurations are colored in gray, with best results highlighted in bold.
IV-C4 Effect of the LKFA and GKFA Modules
We evaluate the contributions of the LKFA and GKFA modules introduced in Section III-D and Section III-E, respectively. As shown in Table VII, removing the LKFA results in a 2.1% drop in the stringent 5°2cm metric, primarily due to the loss of fine-grained geometric features. Furthermore, removing the GKFA leads to a larger decline of 3.0%, indicating its role in preserving global consistency. These results underscore that both local and global geometric information are crucial for establishing accurate correspondences on unseen objects.
Method $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
@@ IV-C4 Effect of the LKFA and GKFA Modules
TABLE VII: Effect of the LKFA and GKFA modules. Default configurations are colored in gray, with best results highlighted in bold.
IV-C5 Effect of the Number of LKFA and GKFA Modules
We investigate the impact of stacking different numbers of keypoint feature aggregation blocks, each containing both a Local Keypoint Feature Aggregator (LKFA) and a Global Keypoint Feature Aggregator (GKFA). As shown in Table VIII, increasing the number of blocks from $S=4$ to $S=12$ consistently improves performance across all metrics, as deeper cascaded aggregation enhances both local and global keypoint representations. However, using $S=16$ blocks leads to performance degradation due to feature redundancy and oversmoothing, which reduce keypoint distinctiveness and impair pose estimation. Overall, $S=12$ achieves the best balance between accuracy and efficiency.
$S$ $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
4 59.6 67.5 76.9 86.3
8 61.4 67.0 79.0 86.8
12 63.2 68.3 80.9 88.0
16 60.7 66.9 79.6 87.6
TABLE VIII: Effect of the number of LKFA and GKFA modules. Default configurations are colored in gray, with best results highlighted in bold.
IV-C6 Effect of Neighborhood Size
To evaluate the influence of neighborhood size $K$ in the LKFA module, we vary the number of nearest neighbors used during local feature aggregation. As shown in Table IX, when $K=1$, the model only uses the center point feature, which limits its ability to capture surrounding geometric and semantic context. Increasing $K$ enhances local structure modeling, with optimal performance achieved at $K=4$. Further increasing $K$ introduces redundant information, leading to no additional gains.
$K$ $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
1 59.6 67.5 77.6 87.0
2 61.1 68.0 79.2 88.0
4 63.2 68.3 80.9 88.0
6 59.7 68.1 77.9 88.4
8 59.9 67.9 77.0 87.1
TABLE IX: Effect of neighborhood size. Default configurations are colored in gray, with best results highlighted in bold.
@@ IV-C7 Effect of Bidirectional Mamba and Feature Sequence Flipping
We also examine the contributions of the bidirectional Mamba architecture and the proposed Feature Sequence Flipping (FSF) strategy within the GKFA module (Section III-E), as summarized in Table X. The results show that replacing the bidirectional Mamba with its unidirectional counterpart (Uni-Mamba) consistently degrades performance, confirming the advantage of bidirectional modeling for global feature aggregation. Additionally, we compare FSF with the commonly used Point Sequence Flipping (PSF) strategy. PSF exacerbates the non-causal nature of point clouds by reversing the sequence along spatial indices, disrupting the alignment between feature order and geometric structure, and leading to performance drops even below the unidirectional baseline. In contrast, FSF preserves the original point order while achieving bidirectional information flow in feature space, effectively improving robustness to the unordered and non-causal characteristics of point clouds.
Method $5^{\circ}2cm$ $5^{\circ}5cm$ $10^{\circ}2cm$ $10^{\circ}5cm$
Uni-Mamba 60.0 66.6 78.5 86.8
@@
TABLE X: Effect of bidirectional Mamba and feature sequence flipping. Default configurations are colored in gray, with best results highlighted in bold.

### v4

@@ IV-B1 Results on REAL275
Table I reports the performance of INKL-Pose compared with state-of-the-art methods on the REAL275 dataset [wang2019normalized]. Specifically, INKL-Pose achieves 84.0% on IoU50, 80.6% on IoU75, 63.2% on 52cm, 68.3% on 55cm, 80.9% on 102cm, and 88.0% on 105cm, demonstrating its strong capability in both pose and size estimation under real-world conditions. Compared with GCE-Pose [li2025gce], the state-of-the-art prior-based method, INKL-Pose achieves significant improvements of 6.2% under the strict 52cm metric. Furthermore, compared to CleanPose [lin2025cleanpose], the current best-performing shape-prior-free method, INKL-Pose yields performance gains of 1.5% under the strict 52cm metric, while introducing only a modest increase in model parameters. In contrast, MK-Pose [yang2025mk] relies on a substantially larger model, yet does not demonstrate commensurate improvements under strict pose metrics.
In addition, compared to our strong baseline AG-Pose [lin2024instance], INKL-Pose achieves notable improvements of 6.2% on the 52cm metric, highlighting the effectiveness of our fine-to-global geometric aggregation strategy. Qualitative comparisons in Fig. 4 further validate the superiority of INKL-Pose. Our method accurately estimates object pose even on camera instances with complex geometric structures, while maintaining a favorable balance between accuracy and model complexity.
@@ IV-B2 Results on CAMERA25
Table II reports the performance of INKL-Pose compared with state-of-the-art methods on the CAMERA25 dataset. As shown in the table, INKL-Pose achieves outstanding accuracy across nearly all evaluation metrics. Although its IoU50 metric reaches 94.4%, slightly lower than SpherePose [ren2025learning] (achieving 94.8%), it significantly outperforms all existing approaches on the remaining metrics, especially under the more rigorous pose accuracy thresholds. Specifically, compared to prior-based methods, INKL-Pose maintains a clear advantage without relying on pre-defined shape priors. In addition, when compared with CleanPose [lin2025cleanpose], the current state-of-the-art shape-prior-free method, INKL-Pose achieves notable gains of 0.7% and 0.8% under the strict 5°2cm and 5°5cm metrics, respectively. These results, along with the strong performance observed on the REAL275 dataset, provide compelling evidence of the robustness and effectiveness of our proposed local-to-global geometric feature aggregation framework. Furthermore, INKL-Pose exhibits strong model efficiency with 16.7M parameters, achieving a fast inference speed of 36 FPS on an NVIDIA RTX 4090D GPU, highlighting its potential for real-time applications.
TABLE III: Quantitative comparisons with state-of-the-art methods on the HouseCat6D Dataset[jung2024housecat6d]. The best results are in bold.
@@
IoU25 IoU50 $^{\circ}2\mathrm{cm} 55$$^{\circ}5\mathrm{cm}$ $10^{\circ}2\mathrm{cm}$ $10^{\circ}5\mathrm{cm}$
GPV-Pose[di2022gpv] 74.9 50.7 3.5 4.6 17.8 22.7
VI-Net[lin2023vi] 80.7 56.4 8.4 10.3 20.5 29.1
SecondPose[chen2024secondpose] 83.7 66.1 11.0 13.4 25.3 35.7
AG-Pose[lin2024instance] 82.4 66.0 11.5 12.6 37.4 42.5
AG-Pose (DINOv2)[lin2024instance] 88.1 76.9 21.3 22.1 51.3 54.3
SpherePose[ren2025learning] 88.8 72.2 19.3 25.9 40.9 55.3
GCE-Pose[li2025gce] - 79.2 24.8 25.7 55.4 58.4
Ours (DINOv2) 90.8 81.9 25.2 27.4 57.5 61.9
@@ IV-B3 Results on HouseCat6D
To further demonstrate the effectiveness of INKL-Pose in handling complex real-world scenarios, we conduct experiments on the more challenging real-world dataset HouseCat6D [jung2024housecat6d]. As reported in Table III, INKL-Pose achieves state-of-the-art performance across all evaluation metrics. Compared with the current state-of-the-art GCE-Pose [li2025gce], our method surpasses it under all evaluation metric, with gains of 2.7% on IoU50, 0.4% on 5°2cm, 1.7% on 5°5cm, 2.1% on 10°2cm, and 3.5% on 10°5cm. Moreover, Fig. 5 shows qualitative pose estimation results under cluttered and heavily occluded conditions, demonstrating that INKL-Pose maintains robust performance in both self-occlusion and inter-object occlusion scenarios.
Figure 4: Qualitative comparisons between INKL-Pose with AG-Pose [lin2024instance] on the REAL275 dataset, where green and red represent the predicted and ground-truth bounding boxes, respectively. Our INKL-Pose demonstrates superior precision in pose estimation, even on camera instances with complex geometries.
Figure 5: Qualitative visualization results of INKL-Pose on the HouseCat6D [jung2024housecat6d] dataset. Green bounding boxes represent the ground truth and red bounding boxes represent the predictions of INKL-Pose. The predictions of INKL-Pose remain highly consistent with the ground truth even under occlusion.
@@ IV-B4 Visualization of Predicted Keypoints
Fig. 6 compares the keypoints predicted by our method with those from the baseline method AG-Pose [lin2024instance] in the NOCS space. To highlight the differences, keypoints are color-coded based on their NOCS errors, linearly mapped from green (error = 0) to red (error = 0.2). As shown in the figure, AG-Pose produces keypoints with higher NOCS errors and uneven surface coverage. Incorporating our keypoint loss functions into AG-Pose leads to moderate improvements in both accuracy and spatial distribution. In contrast, our INKL-Pose achieves the lowest NOCS errors and generates a more uniform and comprehensive keypoint layout that better captures the object’s surface geometry. For example, in the third column showing a mug with a handle, AG-Pose fails to localize the handle, while its variant with our losses partially recovers the region but still exhibits large errors. In comparison, our method accurately delineates the handle contour, resulting in more precise estimations.
@@ IV-C Ablation Studies
In this section, we conduct comprehensive ablation studies to evaluate the contributions of key components proposed in INKL-Pose on the REAL275 [wang2019normalized] dataset.
@@ IV-C1 Effect of the Number of Keypoints
We begin by evaluating how the number of predicted keypoints influences the overall pose estimation performance. As shown in Fig. 7, even with only 16 sparse keypoints, INKL-Pose achieves 56.0% accuracy under the 52cm metric, which is already comparable to many existing state-of-the-art methods. As the number of keypoints increases, the performance steadily improves due to a stronger capacity for capturing fine-grained geometric details, reaching a peak accuracy of 63.2% under the strict 5°2cm metric when using 96 keypoints. Beyond this point, further increasing the number of keypoints results in a slight performance decline, caused by redundancy introduced through overly dense keypoint sampling. Therefore, to achieve a balance between computational efficiency and pose estimation accuracy, we adopt 96 keypoints as the default configuration in our framework.
IV-C2 Effect of the Number of LKFA and GKFA Modules
We investigate the impact of stacking different numbers of keypoint feature aggregation blocks, each containing both a LKFA and a GKFA. As shown in Fig. 8, increasing the number of blocks from $S=4$ to $S=12$ consistently improves performance across all metrics, as deeper cascaded aggregation enhances both local and global keypoint representations. However, using $S=16$ blocks leads to performance degradation due to feature redundancy, which reduce keypoint distinctiveness and impair pose estimation. Overall, $S=12$ achieves the best balance between accuracy and efficiency.
IV-C3 Effect of Neighborhood Size
To evaluate the influence of neighborhood size $K$ in the LKFA module, we vary the number of nearest neighbors used during local feature aggregation. As shown in Fig. 9, when $K=1$, the model only uses the center point feature, which limits its ability to capture surrounding geometric and semantic context. Increasing $K$ enhances local structure modeling, with optimal performance achieved at $K=4$. Further increasing $K$ introduces redundant information, leading to no additional gains.
IV-C4 Effect of Keypoint Supervision
In the Instance-Adaptive Keypoint Detector (IAKD) module (Section III-C), we introduce two auxiliary loss terms, $\mathcal{L}_{surf}$ and $\mathcal{L}_{sep}$, which encourage uniform surface coverage and spatial diversity among keypoints. Table IV summarizes the performance under different supervision strategies. As the table shows, without any keypoint supervision (“None”), the model performs poorly across all metrics, highlighting the necessity of proper guidance. When incorporating the surface coverage loss $\mathcal{L}_{ocd}$ proposed in AG-Pose, the performance further degrades, indicating that this loss is not well suited to our framework. In contrast, the proposed surface loss $\mathcal{L}_{surf}$ significantly improves performance over the unsupervised baseline.
IV-C5 Effect of Positional Embeddings
We further analyze the contribution of positional information to feature representation. In INKL-Pose, positional encodings are introduced at two distinct stages: the embedding $\mathcal{E}_{\mathcal{P}}$ in the Pointwise Feature Encoder (Section III-B), which compensates for the loss of absolute position cues, and $\mathcal{E}_{kpt}$ in the LKFA module (Section III-D), which enhances local spatial awareness. As reported in Table V, removing either embedding individually leads to a noticeable performance drop across all metrics, highlighting their complementary roles in capturing geometric information.
Figure 7: Effect of the number of keypoints.
Figure 8: Effect of the number of LKFA and GKFA modules.
Figure 9: Effect of neighborhood size.
TABLE IV: Effect of keypoint supervision. Default configurations are colored in gray, with best results highlighted in bold.
Method 52cm 55cm 102cm 105cm
w/ ($\mathcal{L}_{sep}$$\mathcal{L}_{surf}$) 63.2 68.3 80.9 88.0
w/ ($\mathcal{L}_{div}$$\mathcal{L}_{ocd}$) 60.9 67.0 79.9 87.2
@@ IV-C5 Effect of Positional Embeddings
TABLE V: Effect of positional embeddings. Default configurations are colored in gray, with best results highlighted in bold.
Method 52cm 55cm 102cm 105cm
@@
IV-C6 Effect of the LKFA and GKFA Modules
We evaluate the contributions of the LKFA and GKFA modules introduced in Section III-D and Section III-E, respectively. As shown in Table VI, removing the LKFA results in a 2.1% drop in the stringent 5°2cm metric, primarily due to the loss of fine-grained geometric features. Furthermore, removing the GKFA leads to a larger decline of 3.0%, indicating its role in preserving global consistency. These results underscore that both local and global geometric information are crucial for establishing accurate correspondences on unseen objects within the same category.
TABLE VI: Effect of the LKFA and GKFA modules. Default configurations are colored in gray, with best results highlighted in bold.
Method 52cm 55cm 102cm 105cm
@@ IV-C7 Effect of Bidirectional Mamba and Feature Sequence Flipping
We also examine the contributions of the bidirectional Mamba architecture and the proposed Feature Sequence Flipping (FSF) strategy within the GKFA module (Section III-E), as summarized in Table VII. The results show that replacing the bidirectional Mamba with its unidirectional counterpart (Uni-Mamba) consistently degrades performance, confirming the advantage of bidirectional modeling for global feature aggregation. Furthermore, we replace the bidirectional Mamba in GKFA with a standard attention-based module consisting of a self-attention layer followed by a feed-forward network (FFN), corresponding to the core sequence modeling component in Transformer architectures. Under identical architectural and training settings, this attention-based variant exhibits a clear performance drop, particularly under the strict $5^{\circ}$2cm metric, and does not outperform the unidirectional Mamba baseline, indicating that Mamba is more effective for capturing long-range dependencies among keypoints in this setting. Additionally, we compare FSF with the commonly used Point Sequence Flipping (PSF) strategy. PSF exacerbates the non-causal nature of point clouds by reversing the sequence along spatial indices, disrupting the alignment between feature order and geometric structure, and leading to performance drops even below the unidirectional baseline. In contrast, FSF preserves the original point order while achieving bidirectional information flow in feature space, effectively improving robustness to the unordered and non-causal characteristics of point clouds.
TABLE VII: Effect of bidirectional Mamba and feature sequence flipping. Default configurations are colored in gray, with best results highlighted in bold.
Method 52cm 55cm 102cm 105cm
w/ Bi-Mamba 63.2 68.3 80.9 88.0
w/ Uni-Mamba 60.0 66.6 78.5 86.8
w/ Attention 59.8 66.8 78.8 86.8
@@
To further analyze the behavior of FSF within the GKFA module, we provide a qualitative visualization of channel-wise feature correlations in Fig. 10. Specifically, Fig. 10(a) shows the cosine similarity matrix of keypoint feature channels before global aggregation, where structured but uneven inter-channel relationships can be observed, reflecting the heterogeneous nature of local geometric and semantic cues. After applying GKFA with bidirectional Mamba and FSF, Fig. 10(b) exhibits a more coherent and structured correlation pattern across feature channels. This suggests that bidirectional aggregation in feature space facilitates more effective information interaction among channels, which is consistent with the quantitative improvements reported in Table VII.
Figure 10: Visualization of cosine similarity matrices of keypoint feature channels. (a) Channel-wise similarity of local keypoint features before GKFA. (b) Channel-wise similarity after GKFA.